OpenStudy (anonymous):

sqrt[3]{125x ^{9}y ^{12}}

OpenStudy (anonymous):

\[\sqrt[3]{125x ^{9}y ^{12}}\]

OpenStudy (anonymous):

You can start by rewriting the pieces, \[e.g. x^9=x^{3+3}=x^3x^3\]

OpenStudy (anonymous):

how did you do that?

OpenStudy (jdoe0001):

check your exponential rules http://www.math-play.com/image-exponents-rules.jpg

OpenStudy (jdoe0001):

\(\large \sqrt[3]{125x ^{9}y ^{12}}\qquad \begin{cases} 5\cdot 5\cdot 5\implies 5^3\implies &125\\ x^3\cdot x^3\cdot x^3\implies x^{3+3+3}\implies &x^9\\ y^3\cdot y^3\cdot y^3\cdot y^3\implies y^{3+3+3+3}\implies& y^{12} \end{cases} \\ \quad \\ \large \sqrt[3]{125x ^{9}y ^{12}}\iff \sqrt[3]{?}\)